OpenStudy (anonymous):

Pam likes to practice dancing while preparing for a math tournament. She spends 80 minutes every day practicing dance and math. To help her concentrate better on math, she spens 20 minutes more dancing than doing math. Part A: Write a pair of linear equations to show the relationship between the number of minutes Pam practices math every day (x) and the number of minutes she practices dances every day (y).

OpenStudy (anonymous):

please help

OpenStudy (mathmale):

I'd suggest you choose variables, such as the usual x and y, to represent the number of minutes Pam spends practicing dance and practicing math, respectively. What is the sum of x and y? Are x and y equal? Why or why not? If they're not equal, how is x related to y?

OpenStudy (anonymous):

Let: d = minutes dancing Let: m = minutes doing math She spends 80 min doing both. d + m = 80 She dances for 20 min more than she does math d = m + 20 You now have a system of equations. Equation #2 is already in d in terms of m, so use that to substitute the expression into equation 1, then solve for the remaining variable. d + m = 80 (m + 20) + m = 80 2m = 60 m = 30 Now that we know "m", we can solve for "d" d = m + 20 d = 30 + 20 d = 50 She dances for 50 min and does math for 30.

OpenStudy (anonymous):

WOW thanks

OpenStudy (anonymous):

@mathmale was silverhotfirefox correct?

OpenStudy (mathmale):

You can check this yourself. What are d and m? Do d and m add up to 80 (as they should)? What does d represent? m? which is the greater of the two? If you take the smaller and add 20, does that give you the larger?

OpenStudy (the_fizicx99):

So creative... https://answers.yahoo.com/question/index?qid=20140115105745AAztHNl

OpenStudy (anonymous):

@SilverHotFireFox in part b: it asks how much time does pam spend on practicing math every day

OpenStudy (anonymous):

wow lol thanks @tHe_FiZiCx99

OpenStudy (anonymous):

part c: is it possible for pam to have spent 60 minutes practicing dance every day?